### v1

# Convexity of the Bergman Kernels on Convex Domains

Abstract.
Let $\Omega$ be a convex domain in $\mathbb{C}^{n}$ and $\varphi$ a convex function on $\Omega$. We prove that $\log{K_{\Omega,\varphi}(z)}$ is a convex function (might be identically $\infty$- ∞) on $\Omega$, where $K_{\Omega,\varphi}$ is the weighted Bergman kernel.
Supported by National Natural Science Foundation of China, No. 12271101
1. Introduction
Let $\Omega$ be a domain in $\mathbb{C}^{n}$ and $\varphi$ a plurisubharmonic (psh) function on $\Omega$. Denote by $A^{2}(\Omega,\varphi)$ the weighted Bergman space of $L^{2}$ integrable holomorphic functions with respect to the weight $e^{-\varphi}$ on $\Omega$, i.e.,
$A^{2}(\Omega,\varphi):=\left\{f\in\mathcal{O}(\Omega):\|f\|_{\Omega,\varphi}^{2}=\int_{\Omega}|f|^{2}e^{-\varphi}<+\infty\right\}.$
The Bergman kernel $K_{\Omega,\varphi}(\zeta,z)$ of $A^{2}(\Omega,\varphi)$ is a function on $\Omega\times\Omega$, which is holomorphic with respect to $(\zeta,\bar{z})$. The function $K_{\Omega,\varphi}(z):=K_{\Omega,\varphi}(z,z)$ is a smooth function on $\Omega$, which is also called the (diagonal) Bergman kernel function. An equivalent characterization of $K_{\Omega,\varphi}(z)$ is the following extremal property:
$K_{\Omega,\varphi}(z)=\sup\left\{|f(z)|^{2}:f\in{A^{2}(\Omega,\varphi)},\ \|f\|_{\Omega,\varphi}\leq 1\right\}.$
In particular, it follows that $\log{K_{\Omega,\varphi}}(z)$ is always a psh function (it can be identically $\infty$- ∞).
The goal of this note is to show the following
Theorem 1.1.
If $\Omega\subset\mathbb{C}^{n}$ is a convex domain, $\varphi$ is a convex function on $\Omega$, then $\log{K_{\Omega,\varphi}(z)}$ is a convex function.
Here and in what follows, we say that a function $\phi:\Omega\rightarrow[-\infty,+\infty)$ is convex if $\Omega$ is a convex domain and
$\phi((1-t)z_{0}+tz_{1})\leq(1-t)\phi(z_{0})+t\phi(z_{1})$
for all $z_{0},z_{1}\in\Omega$ and $t\in[0,1]$. Note that a convex function $\phi$ identically equals to $\infty$- ∞ if $\phi(z_{0})=-\infty$ for some $z_{0}\in\Omega$. Otherwise $\phi$ would be a continuous function on $\Omega$. Thus $\log{K_{\Omega,\varphi}(z)}>-\infty$ if $\Omega$ is bounded and $\varphi(z)>-\infty$. Indeed, it suffices to prove Theorem 1.1 under these additional assumptions, in view of Ramadanov’s theorem.
When $\varphi=0$, we write $K_{\Omega}$ instead of $K_{\Omega,0}$. In the case that $n=1$ and $\Omega\subsetneq\mathbb{C}$ is a convex domain, $K_{\Omega}$ is closely related to the hyperbolic metric $ds_{\Omega}=\lambda_{\Omega}(z)|dz|$ (normalized with Gauss curvature $-1$), via the following identity:
$K_{\Omega}(z)=\frac{1}{4\pi}\lambda_{\Omega}(z)^{2},\ \ \ \forall\,z\in\Omega.$
As a consequence, the convexity of $\log{K_{\Omega}(z)}$ is equivalent to that of $\log\lambda_{\Omega}(z)$. The latter was proved by Caffarelli-Friedman [5] by using PDE techniques, for $u:=\log\lambda_{\Omega}(z)$ is the solution of the Liouville equation
$\Delta{u}=e^{u}.$
Indeed, the convexity of solutions for more general semilinear elliptic PDEs is established in [5].
A complex-analytic proof was given later by Gustafsson [6]. Recall that a conformal mapping $f:\mathbb{D}\rightarrow\Omega$ is called a convex univalent function if $\Omega\subsetneq\mathbb{C}$ is a convex domain. A classical result (cf. Ahlfors [1]) asserts that $f$ is convex univalent if and only if
(1.1) $\mathrm{Re}\,\frac{zf^{\prime\prime}(z)}{f^{\prime}(z)}\geq-1.$
Gustafsson showed that the convexity of $\log\lambda_{\Omega}(z)$ (i.e., the convexity of $\log{K_{\Omega}(z)}$) is equivalent to (1.1), and hence is equivalent to the convexity of $\Omega$.
It is also interesting to point out that the convexity of $\log\lambda_{\Omega}(z)$ has been essentially known earlier. Minda-Wright [9] proved a seemingly stronger result that $-1/\lambda_{\Omega}(z)$ is a convex function if $\Omega\subsetneq\mathbb{C}$ is convex. Since $-\log(-t)$ is a convex and increasing function, it follows that $\log\lambda_{\Omega}=-\log(-(-1/\lambda_{\Omega}))$ is convex on $\Omega$. Their method is a surprisingly simple application of the Schwarz-Pick lemma. Indeed, the convexity of domain $\Omega$ and functions $\log\lambda_{\Omega}(z)$ and $-1/\lambda_{\Omega}(z)$ are all equivalent (see Kim-Minda [7] for more equivalent conditions).
For higher-dimensional cases, some crucial difficulties might arise if one wants to extend the above methods. For example, the Riemann mapping theorem fails in the case of several complex variables, and it is also not known whether $K_{\Omega,\varphi}(z)$ satisfies certain elliptic PDE. Instead, we may reduce the proof of Theorem 1.1 to a celebrated theorem of Berndtsson [2] on subharmonicity properties of Bergman kernels (the case $n=1$ and $\varphi=0$ is due to Maitani-Yamaguchi [8]), by using certain characterization of convexity in terms of subharmonicity (see Lemma 2.1).
We conclude this section with the following question in the light of Minda-Wright [9].
Question .
Let $\Omega\subset\mathbb{C}^{n}$ be a convex domain on $\Omega$. Is $-1/K_{\Omega}(z)^{1/2}$ a convex function?
Remark .
Xu Wang informed the author that one cannot expect the convexity of $-1/K_{\Omega,\varphi}(z)^{1/2}$ for arbitrary convex weight function $\varphi$ even in the case $n=1$. Indeed, take $\Omega=\mathbb{C}$ and $\varphi(z)=\phi(x):=x^{2}$, where $z=x+iy\in\mathbb{C}$. By a result of Berndtsson (cf. [4, §7.1]), we have
$K_{\mathbb{C},\varphi}(z)=\int_{\mathbb{R}}e^{tx-\phi(t)^{2}}=\pi^{1/2}e^{x^{2}}.$
But $-1/K_{\Omega,\varphi}(z)^{-1/2}=-\pi^{-1/4}e^{-x^{2}/2}$ is not convex.
2. Convex functions and subharmonic functions
Let us first consider a smooth function $\varphi$ of one complex variable $t$. The real Hessian of $\varphi$ is the following quadratic form
$\eta\mapsto\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}|\eta|^{2}+\mathrm{Re}\,\left(\frac{\partial^{2}\varphi}{\partial{t}^{2}}\eta^{2}\right),\ \ \ \eta\in\mathbb{C}.$
It is well-known that $\varphi$ is convex if and only if the real Hessian of $\varphi$ is semi-positive. For $\eta\neq 0$, set $\lambda=\eta/|\eta|$. Thus $\varphi$ is convex if and only if
(2.1) $\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}+\mathrm{Re}\,\left(\lambda^{2}\frac{\partial^{2}\varphi}{\partial{t}^{2}}\right)\geq 0$
for all complex number $\lambda$ with $|\lambda|=1$. Since we may always choose $\varphi$ with $\mathrm{Re}\,(\lambda^{2}\partial^{2}\varphi/\partial{t}^{2})\leq 0$, (2.1) implies
$\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}\geq 0,$
i.e., $\varphi$ is a subharmonic function.
Next, let $|\lambda|<1$ and set
$s:=\frac{t-\lambda^{2}\bar{t}}{1-|\lambda|^{4}},$
i.e., $t=t_{\lambda}(s):=s+\lambda^{2}\bar{s}$. Based on (2.1), we have the following criterion for convexity.
Lemma 2.1.
Let $U$ be a convex domain in $\mathbb{C}$ and $V_{\lambda}:=\{s\in\mathbb{C}:t=t_{\lambda}(s)\in{U}\}$. If $\varphi$ is a function on $U$ such that $s\mapsto\varphi(t_{\lambda}(s))$ is subharmonic on $V_{\lambda}$ whenever $|\lambda|<1$, then $\varphi$ is convex on $U$.
Proof.
First of all, we assume that $\varphi$ is smooth. A straightforward calculation yields
$\frac{\partial\varphi(t_{\lambda}(s))}{\partial{s}}=\frac{\partial\varphi}{\partial{t}}\frac{\partial{t_{\lambda}(s)}}{\partial{s}}+\frac{\partial\varphi}{\partial\bar{t}}\frac{\partial\overline{t_{\lambda}(s)}}{\partial{s}}=\frac{\partial\varphi}{\partial{t}}+\bar{\lambda}^{2}\frac{\partial\varphi}{\partial\bar{t}}$
and
$\frac{\partial^{2}\varphi(t_{\lambda}(s))}{\partial{s}\partial\bar{s}}=$ $\,\frac{\partial}{\partial\bar{s}}\left(\frac{\partial\varphi}{\partial{t}}\right)+\bar{\lambda}^{2}\frac{\partial}{\partial\bar{s}}\left(\frac{\partial\varphi}{\partial\bar{t}}\right)$
$=$ $\,\frac{\partial^{2}\varphi}{\partial{t}^{2}}\frac{\partial{t_{\lambda}(s)}}{\partial\bar{s}}+\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}\frac{\partial\overline{t_{\lambda}(s)}}{\partial\bar{s}}+\bar{\lambda}^{2}\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}\frac{\partial{t_{\lambda}(s)}}{\partial\bar{s}}+\bar{\lambda}^{2}\frac{\partial^{2}\varphi}{\partial\bar{t}^{2}}\frac{\partial\overline{t_{\lambda}(s)}}{\partial\bar{s}}$
$=$ $\,(1+|\lambda|^{4})\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}+2\mathrm{Re}\,\left(\lambda^{2}\frac{\partial^{2}\varphi}{\partial{t}^{2}}\right).$
Since $s\mapsto\varphi(t_{\lambda}(s))$ is subharmonic, we have
$(1+|\lambda|^{4})\frac{\partial^{2}\varphi}{\partial{t}\partial\bar{t}}+2\mathrm{Re}\,\left(\lambda^{2}\frac{\partial^{2}\varphi}{\partial{t}^{2}}\right)\geq 0,\ \ \ \forall\,\lambda:|\lambda|<1$
on $U$. Letting $|\lambda|\rightarrow 1$, we obtain (2.1) for any complex number $\lambda$ with $|\lambda|=1$.
In general, we fix an arbitrary convex subdomain $U^{\prime}\subset\subset{U}$. By taking $\lambda=0$, we see that $\varphi$ is a subharmonic function on $U$. In particular, it is locally integrable. Take a cut-off function $\kappa\in{C^{\infty}_{0}(\mathbb{C})}$ with $\kappa\geq 0$ and $\int_{\mathbb{C}}\kappa=1$. Then the convolution
$\varphi_{j}(t)=\int_{\tau\in\mathbb{C}}\varphi(t-j^{-1}\tau)\kappa(\tau)$
is a smooth function defined on $U^{\prime}$ when $j\gg 1$, such that $\varphi_{j}\downarrow\varphi$ as $j\uparrow\infty$. Note that
$\varphi_{j}(t_{\lambda}(s))=\int_{\tau\in\mathbb{C}}\varphi(t_{\lambda}(s)-j^{-1}\tau)\kappa(\tau)=(1-|\lambda|^{4})\int_{\tau^{\prime}\in\mathbb{C}}\varphi(t_{\lambda}(s-j^{-1}\tau^{\prime}))\kappa\circ{t_{\lambda}}(\tau^{\prime}),$
where $\tau^{\prime}=t_{\lambda}^{-1}(\tau)$ and $1-|\lambda|^{4}=\det(t_{\lambda})$ (the determinant of $t_{\lambda}$ as a real linear mapping on $\mathbb{R}^{2}$). It follows that $\varphi_{j}(t_{\lambda}(s))$ is subharmonic with respect to $s\in V^{\prime}_{\lambda}:=\{s\in\mathbb{C}:t_{\lambda}(s)\in{U^{\prime}}\}$. Thus $\varphi_{j}$ is convex on $U^{\prime}$, and so is its decreasing limit $\varphi$. Since $U^{\prime}$ is arbitrary, we complete the proof. ∎
3. Berndtsson’s theorem for convex domains and weight functions
To prove Theorem 1.1, we shall make use of the following modification of Berndtsson’s subharmonicity theorem (cf. [2]).
Theorem 3.1.
Let $\widetilde{\Omega}$ be a bounded convex domain in $\mathbb{C}^{n+1}$ and $\widetilde{\varphi}$ a convex function on $\widetilde{\Omega}$. Set
$\widetilde{\Omega}_{t}:=\left\{z\in\mathbb{C}^{n}:(z,t)\in\widetilde{\Omega}\right\}\ \ \ \text{and}\ \ \ \widetilde{\varphi}_{t}:=\widetilde{\varphi}(\cdot,t).$
Then $t\mapsto\log{K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}(z)}$ is a convex function on $\{t\in\mathbb{C}:(z,t)\in\widetilde{\Omega}\}$.
Proof.
In the case that $\widetilde{\Omega}$ is pseudoconvex and $\widetilde{\varphi}$ is psh, Berndtsson proved that $t\mapsto\log{K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}(z)}$ is a subharmonic function (cf. [2], Theorem 1.1). Take any $\lambda\in\mathbb{C}$ with $|\lambda|<1$ and let $t_{\lambda}$ be given as above. Define a real linear isomorphism $T$ of $\mathbb{C}^{n+1}$ by
$T:(z,s)\mapsto(z,t_{\lambda}(s))=(z,s+\lambda^{2}\bar{s}).$
Set $\widetilde{D}=T^{-1}(\widetilde{\Omega})$, $\widetilde{\psi}=\widetilde{\varphi}\circ{T}$ and
$\widetilde{D}_{s}:=\left\{z\in\mathbb{C}^{n}:(z,s)\in\widetilde{D}\right\}=\widetilde{\Omega}_{t_{\lambda}(s)},\ \ \ \widetilde{\psi}_{s}:=\widetilde{\psi}(\cdot,s)=\widetilde{\varphi}_{t_{\lambda}(s)}.$
It follows that $\widetilde{D}$ is still a convex domain in $\mathbb{C}^{n+1}$ and $\widetilde{\psi}$ is convex with respect to $(z,s)$. In particular, $\widetilde{D}$ is pseudoconvex and $\widetilde{\psi}$ is psh with respect to $(z,s)$, so that Berndtsson’s theorem applies, i.e.,
$s\mapsto\log{K_{\widetilde{D}_{s},\widetilde{\psi}_{s}}(z)}=\log{K_{\widetilde{\Omega}_{t_{\lambda}(s)},\widetilde{\varphi}_{t_{\lambda}(s)}}(z)}$
is a subharmonic function on $\{s\in\mathbb{C}:(z,s)\in\widetilde{D}\}=\{s\in\mathbb{C}:(z,t_{\lambda}(s))\in\widetilde{\Omega}\}$. By Lemma 2.1, $t\mapsto\log{K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}(z)}$ is a convex function on $\{t\in\mathbb{C}:(z,t)\in\widetilde{\Omega}\}$. ∎
4. Proof of Theorem 1.1
Proof of Theorem 1.1.
Given a bounded convex domain $\Omega\subset\mathbb{C}^{n}$ and two points $z_{0},z_{1}\in\Omega$, set
$z_{t}:=(1-t)z_{0}+tz_{1},\ \ \ \forall\,t\in\mathbb{C}$
and $U:=\{t\in\mathbb{C}:z_{t}\in\Omega\}$. Since $U$ is the pre-image of $\Omega$ under the linear mapping $t\mapsto{z_{t}}$, it follows that $U$ is a convex neighbourhood of $[0,1]$. Consider the following domain
$\widetilde{\Omega}:=\{(z,t)\in\mathbb{C}^{n+1}:z+z_{t}\in\Omega,\ t\in{U}\}$
and the function
$\widetilde{\varphi}(z,t):=\varphi(z+z_{t}),\ \ \ (z,t)\in\widetilde{\Omega}.$
It is easy to verify that $\widetilde{\Omega}$ is a convex domain and $\widetilde{\varphi}$ is a convex function on $\widetilde{\Omega}$. By Theorem 3.1, we see that the function $t\mapsto\log{K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}(z)}$ is convex on $\{t\in\mathbb{C}:(z,t)\in\widetilde{\Omega}\}$. Since
$\widetilde{\Omega}_{t}=\Omega-z_{t}:=\{z-z_{t}:z\in\Omega\},$
we have $0\in\widetilde{\Omega}_{t}$ for all $t\in{U}$ and
$K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}(0)=K_{\Omega,\varphi}(z_{t}).$
Thus we obtain the convexity of
$t\mapsto\log{K_{\Omega,\varphi}(z_{t})},\ \ \ t\in{U}.$
In particular, for $t\in[0,1]$,
$\log{K_{\Omega,\varphi}((1-t)z_{0}+tz_{1})}\leq(1-t)\log{K_{\Omega,\varphi}(z_{0})}+t\log{K_{\Omega,\varphi}(z_{1})},$
so that $\log{K_{\Omega,\varphi}(z)}$ is a convex function on $\Omega$. ∎
Remark .
The idea of constructing the domain $\widetilde{\Omega}$ in the proof of Theorem 1.1 goes back to Oka. It is also used by Maitani-Yamaguchi [8] and Berndtsson [2] to prove that $(z,t)\mapsto\log{K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}}(z)$ is jointly psh. In the setting of Theorem 3.1, one can apply the same trick as in [8, 2] to prove the jointly convexity of $(z,t)\mapsto\log{K_{\widetilde{\Omega}_{t},\widetilde{\varphi}_{t}}}(z)$. This is pointed out to the author by Xu Wang.
Acknowledgement
The author would like to thank Bo-Yong Chen for introducing this topic to him together with constant encouragement. He is also grateful to Xu Wang for valuable comments and useful remarks.